the gravitational potential difference between the surface of a planet and a point 20m above is 16J/kg.the work done in moving a 2kg mass by 8m on a slope of 60 degree from the horizontal is?

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Height=8sin60=4√3
Now potential difference for height 4√3=16/20*4√3=16√3/5
Now, WKT,
work done= potential difference *mass
Therefore, work done=16√3/5*2=11.1 J